A train covers a distance between station A and B in 45 minutes. If the speed is reduced by 5 km/hr,it will cover the same distance in 48 min. What is the distance between A&B in km
Answer by mananth(16949)

let the speed of the train be x km/h
distance between AB = 45/60 * x
speed decreased by 5 km/h
speed = (x-5)
d= (x-5) * 48/60 km
But d is same
so
45/60 * x = 48/60*(x-5)
45x=48x-240
3x=240
/3
x= 80
speed of train is 80 km/h
time = 45 minutes = 3/4 hour
d=rt
80*3/4
=60 km
Distance between A & B = 60 km